CAT Chat Minutes

August 25, 2003 

 

General Information

Operation Stats R. Klaffky: 

Delivered beam:  1738.48 Hours
Percentage of scheduled time:  (*) 95.74
Downtime during period:  77.44 Hours
Mean time between faults (MTBF):  41.39 Hours

# 75 08/19 08:00 To 08/21 07:51 47.86

S38 C1 Temp trip [RF]

1.89 Investigation, lost again, bypass, refill

# 81 08/21 09:45 To 08/21 14:06 4.35

IOC S25FB [OPS/CTL]

0.52 Refill

# 82 08/21 14:37 To 08/21 22:16 7.65

Pwr sag ComEd [other]

0.90 Systems recovered - refill

Beamline Downtime Report:  Summary period 8/15-21/2003

Date

Time

Beamline

Description

Downtime

Group

8/16/03

15:30

14-BM

Ion pump controller failure. Temporary replacement provided.

1:10

Beamline

8/17/03

01:55

-

User could not access computer directory; called beamline personnel. Impacted secondary tasks.

1:14

Beamline

8/18/03

10:57

34-ID

Differential pressure on M4 was too low, causing fault.

0:33

ME

8/19/03

08:07

03-ID

Upstream ID not responding to user command.

0:10

XFE

8/19/03

16:42

02-BM

Mode shutter malfunction. Shutter now runs in mono mode only, pending repair during shutdown.

2:00

*

8/19/03

19:30

-

Beamline vacuum failure.

14:59

Beamline

8/21/03

12:50

20-ID

Beam motion inhibiting beamline ability to complete scans.

1:16

CTL

*Pending review by APS management.

   

Machine studies begins tonight at midnight.  Roger gave a summary of the FY 2003 Operational Statistics. Roger will request that CY check on the annual current and get the average.

There will be an APS UofC Review during September 17-19. On September 18th there is an open review of the APS and poster session. Any interested parties are welcome to attend.

Safety J. Alicz:  Jeff Alicz reported that there will be a Machine Shop Training class tomorrow at 10:30 a.m. in the Bldg. 431 Machine Shop.

Jeff also explained why the need for a second signature on some ESAF forms. The second signature is required only on high-risk experiments. If anyone has any questions, contact Jeff or Elroy Chang.

LOM Network Switch Upgrade - Rod Salazar:  Rod reported the upgrade, which will begin September 2nd. Rod also went over changes in the work affecting beamline during the August-September maintenance period.

Work affecting beamlines during the Aug. Sept. Maintenance Period (8-27-03 through 9-26-03)

Information current as of August 25, 2003

Switch Gear Maintenance
X2 - Saturday Sept. 6, 2003
LOM 432 and Sectors 4, 5, 6, and 7
Power will be off from approximately 6 a.m. to 6 p.m. This will also cause e-power bump at ~ 06:30 and at ~17:30
This affects cryo-pumps connected to e-power.
EFOG will contact beamline personnel and establish some 120 V power to their area for necessary items.

Switch Gear Maintenance
X3 - Saturday Sept. 13, 2003
LOM 433 and Sectors 8, 9, 10, and 11
Power will be off from approximately 6 a.m. to 6 p.m. This will also cause e-power bump at ~ 06:30 and at ~17:30
This affects cryo-pumps connected to e-power.
EFOG will contact beamline personnel and establish some 120 V power to their area for necessary items.

Emergency Power Transfer Switch Testing Saturday Sept 6, 2003
Emergency power will be interrupted twice (15 sec. for switch over and then 5 sec. to restore):
Tests affecting Sectors 33 and 35 e-power to cryo-pump at ~ 07:00 to 07:30 and LOM 438
Tests affecting Sector 1 e-power to cryo-pump at ~ 07:30 to 08:00 and LOM 431
Tests affecting Sectors 9, 10, and 11 e-power to cryo-pump at ~ 08:30 to 09:00
Tests affecting Sectors 12 and 13 e-power to cryo-pump at ~ 09:30 to 10:00
Tests affecting Sectors 17, 18, and 19 e-power to cryo-pump at ~ 10:30 to 11:00
Tests affecting Sectors 20 and 22 e-power to cryo-pump at ~ 11:30 to 12:00

NOTE: these are approximate dates

Survey and Alignment
Start date: 9/17/2003 - Sector 14-BM check the alignment of the beamline components for 14BM
Start date: 9/15/2003 - Sector 10-ID align BE window

Front End
Start date 8/27/2003 Sector 10-ID - replace commissioning window assembly, install 1.33 valve on RGA system and bake out, etc.

Start date 8/27/2003 - Sector 18- ID - repair X-BPM blade and replace FEV, bake out, etc.

PSS
Start date: 8/27/2003 Sector 19-ID move 19-ID station D pneumatic door box farther away from the door
8/27/2003 Sector 2-ID upgrade 2-ID station door locks and led indication
8/28/2003 Sector 7-ID pull out four unused love controllers from the PSS DIW
8/27/2003 Sectors 8-35 pulizzi plug-in configuration change: a) in sectors 8-35, unplug pulizzi power controllers located in the PSS racks from the UPSs and plug them into the blue receptacles powered by the central UPS; b) follow the PSS task workflow checklist to clear all faults; c) notify Roy Emerson about the power cycling that took place.
8/27/2003 Sectors 1-ID and 2-ID pull out obsolete PSS DIW cables.
8/27/2003 Sectors 3-ID ground 2 PSS control panels (upper 15Us), for stations A/B and D.
8/27/2003 Sector 2-ID ground 4 PSS control panels (upper 15Us), for stations A, B, D and E.
8/27/2003 Sector 23-ID FE shutter work; disconnect PSS components from PS2 and safety shutters on 8/27/2003, reinstall PSS switch assemblies when PS2, SS1 and SS2 are ready (~9/10/03), calibrate PSS switches and shutter motion, and validate PSS wiring and controls.
8/26/2003 Sector 4-ID BLEPS work; re-route field device cables to the new rack #2 location and terminate. Move Panel View #2 (PV2) to its new location. Pull the blue hose communication cable from rack #1 to the remote I/O adapter in rack #2, then to PV2. Perform test procedures.
8/27/2003 Sectors 3, 4, and 5 pull new PSS DIW wiring, install new LOVE Controller chassis, perform DIW validations.
8/27/2003 3-ID check SS1 position switches and other parts
8/27/2003 - 14-BM-C relocate PSS search Box 1

PSS Re-validation Schedule: 

8/27-28

10ID

8/27-29

2ID

9/3-5

5BM

9/3-5

6ID

9/8-10

8ID

9/8-10

9BM

9/15-17

12ID

9/15-17

34ID

9/22-24

35ID

9/22-23

35BM

9/29-9/30

4ID

Low-Temperature Hot Water Lines - This work will begin on August 27th and continue through September 26th. The area at column 98 will be drained on August 27th where a gasket needs to be replaced or repaired. This should take only one or two days. This will affect LOM 433 and the experimental hall floor in that area. The affect will be loss of heating, so we may experience some cool temperatures if the weather does not cooperate. As soon as this repair is made, heat will be restored. Insulation of piping will begin no sooner than August 27th. A schedule of areas will be given next week. The work will require moving apparatus on the floor so that workers can access the pipe joints. EFOG will work with beamline personnel to move (as required) and protect apparatus on the floor. Schedule of re-insulating is forthcoming.

The tentative schedule is:
Sectors 1, 2, 3, and 4 will be worked on during the period of Aug. 27th to Sept. 5th working clockwise starting at sector 1

Sectors 35, 34, 33, and 32 will be worked on during the period of Aug. 27th to Sept. 5th working counter clockwise starting at sector 35

Sectors 5, 6, 7, and 8 will be worked on during the period of Sept. 8th to Sept. 12th working clockwise starting at sector 5

Sectors 23, 22, 21, and 20, will be worked on during the period of Sept. 8th to Sept. 12th working counter clockwise starting at sector 23

Sectors 9, 10, 11, and 12 will be worked on during the period Sept. 15th to Sept. 19th working clockwise starting at sector 9

Sectors 19, 18, 17, and 16 will be worked on during the period Sept. 15th to Sept. 19th working counter clockwise starting at sector 19

Sectors 13, 14, and 15 will be worked on during the period Sept. 22nd to Sept. 26th working clockwise starting at sector 13

UPS System on SR Mezzanine
8/28/2003 Emergency power will be shut down for one to two hours beginning with Sector 17, 18, 19 and 20 e-power to cryo-pump at ~ 07:30 to 09:30 LOM 435 (Hoods), card key system, PA system, emergency lighting in LOM 435.  Cold Rooms in sectors 17 and 19.

Process Water
Work in sectors 3,4, and 8 during maintenance period.

Sector 34 Install new throttle valve for M4 - 7/28/03

ID Work
Sectors 3US.3DS (remove, re-measure, reinstall, calibration and maintenance 8/27/2003 to 9/19/2003)
Sector 4DS remove and replace, calibration and maintenance 8/27/2003 to 9/19/2003)
Sector 10-ID remove, measure and reinstall (8/27/2003 to 9/19/2003)
Sector 11-ID install new encoder

Shutter Work
Sector 2-BM (mode shutter) work will begin near end of maintenance period
Sector 16-ID-B located in A enclosure. Start 8/27/03 - reinstall near end of maintenance period.
If you require cryo-pump maintenance, and/or enclosure door maintenance during this period, please contact your Floor Coordinator to make arrangements.

Sector 5 PSS work will not be done during this shutdown; will be scheduled for the next shutdown.